# Probabilistic Kernel Function for Fast Angle Testing

### v1

@@ B.3 Algorithms Related to KS1 and KS2
$\bm{q}$ is the query, $k$ denotes the value of top-$k$, $G$ is the graph index
@@
[…]
@@
17             if KS2_Test($\bm{w},\bm{v},\bm{q},\delta$) = true (see ineq. (10)) then
@@
B.4 Numerical Computation of Reference angle
@@
Appendix C Additional Experiments
C.1 Datasets and Parameter Settings
@@
(1) CEOs: The number of projection vectors, $m$, was set to 2048, following the standard setting in the original paper [25].
(2) KS1: For KS1, $L$ was fixed to 1, as multiple levels are not necessary for angle comparison. The number of projection vectors, $m$, was also set to 2048, consistent with CEOs. We evaluated both KS1($S=S_{\text{sym}}(2048,1)$) and KS1($S=S_{\text{pol}}(2048,1)$).
@@
(5) HNSW+PEOs: Following the suggestions in [22], we set $L$ to 8, 10, 15, 15, 16, and 20 for the six real datasets, sorted in ascending order of dimension. Additionally, $\epsilon$ was set to 0.2, and $m=256$ to ensure that each vector ID could be encoded with a single byte.
@@
C.2 KS1 vs. CEOs under Different Settings
@@
C.3 ANNS Results under Different $k$’s
@@
C.4 The Impact of $L$ on the Other Datasets
@@
C.5 Results of NSSG+KS2
We also implement KS2 on another state-of-the-art similarity graph, NSSG [13]. The parameter settings for NSSG are the same as those in [22]. From the results in Fig. 8, we observe that NSSG+KS2 outperforms NSSG+PEOs on all datasets except for GIST, which indicates that the superiority of KS2 is independent of the underlying graph structure.

### v2

@@ C.4 Algorithms Related to KS1 and KS2
$\bm{q}$ is the query, $k$ denotes the value of top-$k$, $G$ is the similarity graph
@@
/* an ordered list of results, $|R|\leq efs$ $P\leftarrow\{\,$ entry node $v_{0}\in G\,\}$ ;
@@
5    if KS2_Test($\bm{w},\bm{v},\bm{q},\delta$) = true (Ineq. (9)) then
@@
C.5 Numerical Computation of Reference angle
@@
Appendix D Additional Experiments
D.1 Datasets and Parameter Settings
@@
(1) CEOs: The number of projection vectors, $m$, was set to 2048, following the standard setting in Pham (2021).
[…]
@@
(5) HNSW+PEOs: Following the suggestions in Lu et al. (2024), we set $L$ to 8, 10, 15, 15, 16, and 20 for the six real datasets, sorted in ascending order of dimension. Additionally, $\epsilon$ was set to 0.2, and $m=256$ to ensure that each vector ID could be encoded with a single byte.
@@
D.2 KS1 vs. CEOs under Different Settings
@@
D.3 ANNS Results under Different $k$’s
@@
D.4 The Impact of $L$ on the Other Datasets
@@
D.5 Results of NSSG+KS2
We also implement KS2 on another state-of-the-art similarity graph, NSSG (Fu et al., 2022). The parameter settings for NSSG are the same as those in Lu et al. (2024). From the results in Fig. 8, we observe that NSSG+KS2 outperforms NSSG+PEOs on all datasets except for GIST, which indicates that the superiority of KS2 is independent of the underlying graph structure.
@@ D.5 Results of NSSG+KS2
Statements on the Use of Large Language Models
We used LLMs to polish writing only. We are responsible for all the materials presented in this work.